OpenStudy (buttermequeasy):

An architect needs to determine the slope between two points on a ski lift. The two points have been identified as (15, 45) and (150, 60), where x is the horizontal distance and y is the vertical distance from the bottom of the lift. Assuming the lift runs in a straight line, what is the slope of the line between the two points? Help please?

OpenStudy (mertsj):

Do you know the slope formula?

OpenStudy (buttermequeasy):

Yes, I do. It's y=mx+b

OpenStudy (mertsj):

That is the slope-intercept form of a linear equation. The slope formula is y2-y1 over x2-x1 Subtract the y values. Subtract the x values. Make a fraction with the first number on top.

OpenStudy (mertsj):

So in this case: 60-45/150-15 or 15/135 = 1/9 and that is the slope

OpenStudy (buttermequeasy):

Oh.

OpenStudy (mertsj):

Do you understand?

OpenStudy (buttermequeasy):

Not really...